What is the equilibrium constant K for a reaction which has ΔrG° = 24.0 kJ mol-1 , valid at a temperature of 36.7 °C?
Answer #1

Equilibrium constant, K = 8.99 × 10-5​​​​​​

Relation between ArGo and K AG=-RT hk where is : T= temperature in K R= 8.314 J/molk T= 36.7°C = (367+ 273:15) K = 309.85k No